Determine whether the following series converge, converge absolutely, converge conditionally, or diverge.

  1. "\\displaystyle\\sum_{k=1}^\t\u221e" "\u221ak\/k^2+k+1"
  2. "\\displaystyle\\sum_{k=1}^\t\u221e" "k!\/(2^k*k^4)"
  3. "\\displaystyle\\sum_{k=1}^\t\u221e" "[k+(1\/k)]^2\/(k^2+1)"3/2

( As I speak English as a Second language, Can you kindly explain a little bit so that I can do other similar problems by myself and any resources where I should study further theories of Sequence and Series) Thank you.


1
Expert's answer
2021-10-19T01:53:50-0400
  1. Consider the given series: "\sum_{k=1}^{\infty }\frac{\sqrt{k}}{k^2+k+1}"


Use Limit Comparison Test, to test the convergence of the give series.

Let "a_{k}=\frac{\sqrt{k}}{k^2+k+1}"

Take "b_{k}=\frac{\sqrt{k}}{k^2}=\frac{1}{k^{\frac{3}{2}}}"

Using Limit Comparison Test, we have

"\lim_{k\rightarrow \infty }\frac{a_{k}}{b_k}=\lim_{k\rightarrow \infty }\frac{\sqrt{k}}{k^2+k+1}\times \frac{k\sqrt{k}}{1}=\lim_{k\rightarrow \infty }\frac{k^2}{k^2+k+1}=1\neq 0"

So, Limit Comparison Test can be applied

And the series "\sum_{k=1}^{\infty }\frac{1}{k^{3/2}}" is convergent by using "p-" series "(p=\frac{3}{2}>1)"

So, By Limit Comparison Test, the given series "\sum_{k=1}^{\infty }\frac{\sqrt{k}}{k^2+k+1}" is also convergent.

Observe that the given series is a series of positive terms.

Therefore, the given series is absolutely convergent.

2.Consider the series: "\sum_{k=1}^{\infty }\frac{k!}{2^{k}\times k^{4}}"

Use Ratio Test to test the convergence of the given series.

Let "a_{k}=\frac{k!}{2^{k}\times k^{4}}"

Then, "a_{k+1}=\frac{(k+1)!}{2^{k+1}\times (k+1)^{4}}"

By using Ratio Test, we have "L=\lim_{k\rightarrow \infty }\frac{a_{k+1}}{a_{k}}"

Then, we get "L=\lim_{k\rightarrow \infty }\frac{(k+1)!}{2^{k+1}\times (k+1)^{4}}\times \frac{2^{k}\times k^{4}}{k!}"

By simplification, we get

"L=\lim_{k\rightarrow \infty }\frac{k+1}{2}\times (\frac{k}{k+1})^{4}=\frac{1}{2}\times (1)\times (\infty )=\infty >1"

So, the given series is divergent by Ratio Test.

Therefore, the given series is divergent by Ratio Test.

3.Consider the series: "\sum_{k=1}^{\infty }\frac{[k+(\frac{1}{k})]^2}{(k^{2}+1)^{3/2}}"

Let "a_{k}=\frac{(k^2+1)^2}{k^2(k^2+1)^{3/2}}"

Use Limit Comparison Test to test the convergence of the given series

Take "b_{k}=\frac{k^4}{k^5}=\frac{1}{k}"


By using Limit Comparison Test, we have

"L=\lim_{k\rightarrow \infty }\frac{a_{k}}{b_{k}}=\lim_{k\rightarrow \infty }\frac{(k^2+1)^2}{k^2(k^2+1)^{3/2}}\times \frac{k}{1}"

Then, we get "L=1\neq 0"

So, Limit Comparison Test can be applied.

Now "\sum_{k=1}^{\infty }b_k=\sum_{k=1}^{\infty }\frac{1}{k}" is divergent by using "p-" series "(p=1)"


Therefore, by Limit Comparison Test, the given series "\sum_{k=1}^{\infty }\frac{[k+(\frac{1}{k})]^2}{(k^{2}+1)^{3/2}}" is also divergent.